OpenStudy (anonymous):

The voltage V in volts produced by a battery is a linear function of the i in amperes drawn from it .The opposite of the slope of the line repesents the battery 's effective resistance R in ohms .For a certain battery , V=12.0 when i = 10.0 and V=8.4 when i=10.0 a) What is the effective resistance of the battery ? b) Find the voltage that the battery would producfe when the current is 25.0 amperes The key for: a=0.4o hm ,b = 2.4 volts now a)= \frac{12-8.4}{1.0-10.0} \frac{3.6}{9=0.4ohm I can't fig b)

OpenStudy (dumbcow):

V = -0.4i + b 12 = -.4(1) +b 12.4 = b V = -.4i + 12.4 i = 25, solve for V V = -.4(25) +12.4 V = 2.4

OpenStudy (anonymous):

look like the key is wrong?

OpenStudy (anonymous):

what answr a ?

OpenStudy (dumbcow):

how so?

OpenStudy (anonymous):

the key for a is 0.4 ohm

OpenStudy (dumbcow):

slope is negative, resistance is opposite of slope so positive 0.4

OpenStudy (anonymous):

\[\frac{12-8.4}{1.0-10.0}=\frac{3.6}{9} = 0.4 ohm

OpenStudy (anonymous):

match with the key for a , but b can match 2.4